OpenStudy (anonymous):

I'm very confused... Find the inverse Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)=x+a/divided by b and g(x)=cx-d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (anonymous):

@triciaal please help me

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@zepdrix

OpenStudy (danjs):

The inverse of a function, if there is one, is a flip of the function across the line y = x

OpenStudy (triciaal):

still answering your other question!

OpenStudy (anonymous):

so an inverse would be y=x+3 x=y-3

OpenStudy (anonymous):

this is the only question i asked? i made an account because i was so stumped on this question

OpenStudy (triciaal):

you don't flip the operation sign just the variables

OpenStudy (anonymous):

so x=y+3 and y=x+3

OpenStudy (triciaal):

@need_algebra_help sorry

OpenStudy (anonymous):

?

OpenStudy (triciaal):

I am sorry. the other question was next to yours.

OpenStudy (danjs):

the two functions finding the constants, shows that the inverse of + is -, the inverse of multiply is divide... maybe

OpenStudy (anonymous):

alright i now understand what an inverse is but i can't apply it to this question

OpenStudy (danjs):

find the inverse of f(x), compare that to the g(x) function, some of the constants, a,b,c,d need to be the same maybe,

OpenStudy (triciaal):